# Scale-invariant attention

### v1

@@ Appendix H The optimal lengthscale, $\tau$, is around 10
By introducing scale-invariant attention, we introduce one extra lengthscale hyperparameter, $\tau$, which represents the size of the ‘chunks’ we attend over. To select $\tau$, we trained at 4k for ${\tau\in\{10^{-2},10^{-1},10^{0},10^{1},10^{2}\}}$, and compared validation losses (shown in Fig. 5). Validation performance is very similar amongst different $\tau$ at the training context length (though $\tau=10$ is strictly best), but as we extend to out of distribution context lengths (64k, 16$\times$ the training context length) the benefit of $\tau=10$ becomes clearer.
@@ I.1 Entropy scaling of regular attention
Figure 6: Scaling of expected entropy-in-range for standard attention with an independent Gaussian assumption on the logits. We calculate expected entropies in the range $[t,t\Delta]$ for different $t$ and $\Delta\in\{2,5,10\}$, with the lengthscale $\tau$ set to $10$. The dashed lines show the best linear fit of the data. We empirically find that standard Gaussian logits give logarithmic entropy (left panel).
@@ I.3 Alternative scale-invariant attention variants
In our analysis in Section 3.2 we assume that the unmodified logits (query-key products), $\bar{L}_{t}$’s, are standard Gaussian. In this section, we empirically verify the validity of this assumption.
We consider a Llama-3B model (Grattafiori et al., 2024), with the introduction paragraph of a Wikipedia page  [^1] https://en.wikipedia.org/wiki/New_England as the input, which contains 454 tokens. In Fig. 9 we show quantiles of $\bar{L}_{t}$’s for each layer, aggregated over the input and the attention heads in each layer. However, we do not aggregate over the beginning of sequence token, as it is an outlier attention sink (Gu et al., 2024).
Figure 9: Quantile-Quantile (QQ) plots of attention logits in a Llama-3B model (blue line), with theoretical quantiles of a Gaussian shown by the red line.

### v2

@@ Appendix H The optimal lengthscale, $\tau$, is around 10
[…]
@@ I.1 Entropy scaling of regular attention
Figure 6: Scaling of expected entropy-in-range for standard attention with an independent Gaussian assumption on the logits. We calculate expected entropies in the range $t,t\Delta]$ for different $t$ and $\Delta\in\{2,5,10\}$, with the lengthscale $\tau$ set to 1010. The dashed lines show the best linear fit of the data. We empirically find that standard Gaussian logits give logarithmic entropy (left panel).
@@ I.3 Alternative scale-invariant attention variants
I.4 Infini-Attention
During the review process, we were asked to compare scale-invariant attention to Infini-attention (munkhdalai2024leave). Since infini-attention is a method for compressing the KV-cache, it is slightly different to the other methods we compare to (which primarily control entropy), and so we include these results separately. See Table 3, which shows that while infini-attention gives long-context generalization, it is not as strong as our method.
Table 3: Final mean validation losses ($\pm 1$ standard error across 3 seeds) for different methods when training with 4k context length on a 162M parameter GPT-2-style model.
Method Val @ 4k Val @ 16k Val @ 64k
RoPE 3.261 $\pm$ 0.001 3.936 $\pm$ 0.010 5.260 $\pm$ 0.014
Infini-RoPE 3.296 $\pm$ 0.003 3.302 $\pm$ 0.004 3.310 $\pm$ 0.008
Infini-pp-RoPE 3.295 $\pm$ 0.003 3.303 $\pm$ 0.005 3.311 $\pm$ 0.009
I.5 Larger models
We also investigated the ability for scale-invariant attention to generalise to longer contexts in larger models by continual pretraining Llama 2 7B (touvron2023llama). We chose Llama 2 for this experiment because it is one of the few models that have not already mid-trained at a longer context length.
Specifically, we continually pretrained at 4k context length after replacing the default attention mechanism with various other methods, and we looked at validation loss at out-of-distribution lengths (4k/16k/64k). We trained using the  torchtune library, using data from FineWeb (penedo2024fineweb), with AdamW and a learning rate of $2\times 10^{-5}$.
We see in Table 4 that changing the attention mechanism degraded the loss at the training context length, but the performance of scale-invariant $p$-RoPE far exceeds the other methods at 16k and 64k.
Table 4: Validation performance when fine-tuning Llama-2 7B with different attention methods on $\sim$50M tokens. RoPE and RoPE+NTK (denoted *) were not fine-tuned.
Method Val loss @4k Val loss @16k Val loss @64k
RoPE* 1.968 7.036 8.815
pp-RoPE 2.029 3.504 6.800
NoPE 4.152 6.172 7.730
RoPE+NTK* 1.988 3.090 7.523
YaRN 1.957 7.004 8.763
LogN+RoPE 1.966 6.932 8.726
LogN+pp-RoPE 2.049 2.984 6.224
LogN+NTK 1.966 3.063 7.413
ALiBi 2.750 2.745 2.744
Scale-invariant $p$-RoPE (ours) 2.163 2.193 2.252
@@ I.5 Larger models
In our analysis in Section 3.2 we assume that the unmodified logits (query-key products), $\bar{L}_{t}$’s, are standard Gaussian. In this section, we empirically verify that the logits are Gaussian by looking at QQ-plots.
We consider several sizes of model, including 1B and 8B Llama variants (grattafiori2024llama), and Gemma 2 27B team2024gemma. We calculate logits with the introduction paragraph of a Wikipedia page  [^1] https://en.wikipedia.org/wiki/New_England as the input. In Figs. 9,10,11,12,13,14 we show quantiles of $\{\bar{L}_{t}\}_{t>0}$’s (i.e. lower triangular part of the $QK^{T}$ matrix) for each layer, aggregated over the input and the attention heads in each layer. Note that we do not aggregate over the ‘beginning of sequence’ token, as it is an outlier attention sink (gu2024attention).
Figure 9: Quantile-Quantile ($Q$) plots of attention logits (without RoPE applied) in a Llama-1B model (blue line), with theoretical quantiles of a Gaussian shown by the red line.
Figure 10: Quantile-Quantile ($Q$) plots of attention logits (with RoPE applied) in a Llama-1B model (blue line), with theoretical quantiles of a Gaussian shown by the red line.
Figure 11: Quantile-Quantile ($Q$) plots of attention logits (without RoPE applied) in a Llama-8B model (blue line), with theoretical quantiles of a Gaussian shown by the red line.
Figure 12: Quantile-Quantile ($Q$) plots of attention logits (with RoPE applied) in a Llama-8B model (blue line), with theoretical quantiles of a Gaussian shown by the red line.
Figure 13: Quantile-Quantile ($Q$) plots of attention logits (without RoPE applied) in a Gemma 2 27B model (blue line), with theoretical quantiles of a Gaussian shown by the red line.
Figure 14: Quantile-Quantile ($Q$) plots of attention logits (with RoPE applied) in a Gemma 2 27B model (blue line), with theoretical quantiles of a Gaussian shown by the red line.